OpenStudy (anonymous):

What is the exact value of the expression the square root of 7. + 2the square root of 6. + 3the square root of 7.? the square root of 20. 6 the square root of 20. 3 the square root of 7. + 2 the square root of 6. 4 the square root of 7. + 2 the square root of 6.

OpenStudy (anonymous):

\[\sqrt{7} + \sqrt[2]{6} + \sqrt[3]{7}\]